Directions. Solve the following problems. Show your solutions.


I. For the voltaic cell, 2 Na(s) + Fe2+(aq) → 2 Na+


(aq) + Fe(s)


a) Write the two half-reactions.


b) Write the cell notation.


c) Calculate E ° for the cell.


II. Consider the fission reaction 1


0n + 235


92U →


89


37Rb + 3 0


-1e + 3 1


0n + 144


58Ce


How much energy in kilojoules is released per gram of Uranium- 235?


III. For Al-28, calculate the binding energy in kilojoules /mol?


IV. Sketch a portion of the acrylonitrile polymer ( Refer to Table 23.1) assuming it is a


a) head-to-tail polymer


b) head-to-tail, tail-to-tail polymer

1
Expert's answer
2022-02-06T11:19:11-0500

To determine the E? value of a cell, we determine the anode and cathode half cell reactions and then calculate the E? by the help of the formula:


E ? ( cell ) = E? (cathode) - E? (anode)


The half-cell reactions are written as below:


Anode: Na(s) → Na+ (aq) + e- E? (anode) = -2.714V


Cathode: Fe2+ (aq) + 2e- → Fe(s) E? (cathode) = -0.409V


So, E? (cell) = -0.409 – (-2.714) = 2.305V